OpenStudy (anonymous):

Quotient Rule: z^2+1/Z^(1/2) Their answer is: Z^(1/2)(3-z^(-2)/(2) I can't do the simplification! Yes z and Z was a typo only 1 variable here

OpenStudy (anonymous):

When in doubt, Product Rule: \[\huge z^2+z^{-\frac{1}{2}}\]

OpenStudy (anonymous):

but if after derivative you've gotten: \[\huge2z-\frac{1}{z^{\frac{3}{2}}}\] the simplification involves factoring out a 1/2z to make the bottom look nice: \[\huge 2z^{\frac{1}{2}}(z^{\frac{1}{2}})-\frac{z^{\frac{1}{2}}}{z^2}\] \[\huge z^{\frac{1}{2}}(2z^{\frac{1}{2}}-\frac{1}{z^2})\] ....hmmm i can't get their answer either

OpenStudy (amistre64):

assuming z and Z are just typos and that they are actually the same variable: (z^2+1)*z^(-1/2) z^(3/2)+ z^(-1/2) 3/2 z^(1/2)-1/2 z^(-3/2) (3z^(2)-1)/2z^(3/2) z^2(3-z^(-2))/2z^(3/2) z^(1/2) (3-z^(-2))/2

OpenStudy (amistre64):

im not sure if the "answer" is that, it seems to be an oversimplification to me that borders on complicating it.